OpenStudy (anonymous):

why can we not multiply by 2pi/360 when converting to raidians?

OpenStudy (anonymous):

\[180 degrees=\pi radians.\] \[or 360 degrees= 2 \pi radians.\] \[1 degree=\frac{ 2 \pi }{ 360 }\] for x degree=?

OpenStudy (anonymous):

i was converting 2pi/3 to degrees and when i multiply it by 360/2pi i get like 1184.23423 but i know it equals 45 degrees, i was wondering why this is

OpenStudy (anonymous):

well actually it equals 120 but i am still confused on why my calculator says its over 1000

OpenStudy (anonymous):

You multiplied a radian by a radian.

OpenStudy (anonymous):

Maybe you typed something in wrong or the calculator memory wasn't cleared out.

OpenStudy (anonymous):

should i put it in degree when doing this?

OpenStudy (anonymous):

I just used both formats and got the same answer (120).

OpenStudy (anonymous):

i get the right answer when i use 180/pi though. idk my calculator may need to be cleared then

OpenStudy (anonymous):

It's also possible that you accidentally hit a wrong button a few times. Those buttons are pretty small.

OpenStudy (anonymous):

i have no idea.... thank you for the help though, idk why 180/pi works but not 360/2pi

OpenStudy (anonymous):

It should work. There is no reason for that not to since they are equivalent fractions. Maybe you have a ghost in your calculator. :-)